OpenStudy (anonymous):

Find the domain and range of: a. (1,2),(2,-4),(3,-10),(4,-16) Domain {1,2,3,4} Range {2,-4,-10, 16} can someone show full work for how this is found?

OpenStudy (anonymous):

Domain {1,2,3,4} Range {2,-4,-10, 16} Please ask me if you did not get it?

OpenStudy (cwrw238):

the domain is the set offirst values in the brackets and the range the set of second values

OpenStudy (cwrw238):

yw

OpenStudy (cwrw238):

the domain i the possible values of x which is the set of all real numbers range is the same

OpenStudy (cwrw238):

the range is all possible values of f(x)

OpenStudy (cwrw238):

ok?

OpenStudy (cwrw238):

the inverse function of (a) is found by reversing each pair of numbers that is (2,1), etc

OpenStudy (cwrw238):

so domain is {2,-4,-10,-16} and range is {1,2,3,4}

OpenStudy (cwrw238):

(b) to find inverse make x the subject of the formula f(x) = 3x - 1 3x = f(x) + 1 x = (f(x) + 1) / 3 now replace f(x) by x inverse = f-1(x) = (x + 1) / 3

OpenStudy (cwrw238):

domain of the inverse is all real numbers range is all real numbers

OpenStudy (cwrw238):

:) is that a smile? yw

OpenStudy (cwrw238):

btw where do i get a list of these symbols?